OpenStudy (trojanpoem):

integrate : 1/[(x^2-2)sqrt(x^2-1)] dx

MsBrains (ms-brains):

I think this might be able to help. c; http://prntscr.com/apfe8g

OpenStudy (irishboy123):

this doesn't give the same answer as Wolfram but I think it looks ok-ish \(I = \int ~ \dfrac{1}{(x^2 - 2) \sqrt {x^2 -1}} ~ dx\) \(x = \cosh u, ~ dx = \sinh u ~ du\) so \(\int \dfrac{1}{(\cosh^2 u - 2) \sqrt {\cosh^2 u -1}} ~ \sinh u ~ du \) \(= \int \dfrac{1}{(\cosh^2 u - 2) \sinh u} ~ \sinh u ~ du \) \(= \int \dfrac{1}{(\sinh^2 u - 1) } ~ du \) then \(\sec p = \sinh u, ~ \sec p \tan p ~ dp = \cosh u ~ du, ~ du = \dfrac{\sec p \tan p }{\cosh u}~ dp\) and \(\cosh u = \sqrt{\sec^2 p + 1}\) \(\Rightarrow \int \dfrac{1}{(\sec^2 p - 1) } ~ \dfrac{\sec p \tan p}{\sqrt{\sec^2 p + 1}} ~ dp \) \(\Rightarrow \int \dfrac{1}{\tan^2 p } ~ \dfrac{\sec p \tan p}{\sqrt{\sec^2 p + 1}} ~ dp \) \(\Rightarrow \int \dfrac{1}{\frac{\sin p}{\cos p} } ~ \dfrac{\frac{1}{\cos p} }{\sqrt{(\frac{1}{\cos p})^2 + 1}} ~ dp \) \(\Rightarrow \int ~ \dfrac{1 }{\sqrt{\tan^2 p + 1}} ~ dp \) \(\Rightarrow \int ~ \cos p ~ dp \) \( = \sin p + C\) !!! from \(\sec p = \sinh u\) \(\sin p = \sqrt{1 - cosech^2 u} \) So, \(I = \sqrt{1 - cosech^2 \cosh^{-1} x} ~ + C \)

OpenStudy (trojanpoem):

@Ms-Brains , Not sure how to take it from there. I tried to start it by letting u = x^2 - 1 ending up at a stalemate. @IrishBoy123, I think this step should be ? (excuse me, If I am mistaken) \[\int\limits_{}^{}\frac{ 1 }{ \sin p \sqrt{(\frac{ 1 }{ \cos^2x}) + 1}}\] \[\int\limits_{}^{}\frac{ 1 }{ \sqrt{\tan^2x + \sin^2p}}\]

OpenStudy (trojanpoem):

* replace the x with p

OpenStudy (irishboy123):

you're quite right! thank you.

OpenStudy (trojanpoem):

\[u^2 = x^2 - 1\] \[dx = \frac{ u du }{ \sqrt{u^2 +1} }\] \[\frac{ \frac{ udu }{ \sqrt{u^2 + 1}} }{ (u^2 - 1) u}\] \[\int\limits_{}^{}\frac{ du }{ (u^2 - 1)\sqrt{u^2 + 1} }\] let u = tantheta \[u = \tan \theta , du = \sec^2 \theta\] \[\int\limits_{}^{} \frac{ \sec^2 \theta d \theta }{ (\tan^2 \theta - 1) \sec \theta }\] \[\int\limits_{}^{}\frac{ \frac{ 1 }{ \cos \theta } d \theta }{ \frac{ \sin^2 \theta }{ \cos ^2 \theta } - 1 }\] \[\int\limits_{}^{}\frac{ \cos \theta d \theta }{ -(\cos^2 \theta - \sin^2 \theta) } = \int\limits_{}^{} \frac{ d \sin \theta }{ 2 \sin^2 \theta - 1 }\] \[\frac{ 1 }{ \sqrt(2) } \int\limits_{}^{} \frac{ d[\sqrt{2} \sin \theta] }{ [\sqrt{2}\sin \theta]^2 - 1}\] \[\frac{ -1 }{ \sqrt{2} } \tanh^-1(\sqrt{2}\sin \theta) + c\] \[\frac{ -1 }{ \sqrt{2} } \tanh^{-1}(\sqrt(2) \frac{ \sqrt{x^2 - 1} }{ x }) + c\]

OpenStudy (irishboy123):

cool!